OpenStudy (anonymous):

solve for n algebraically. (2n+1) chose 2 = n+50

OpenStudy (mathmate):

(2n+1)choose 2 = (2n+1)!/((2n-1)2!)=n(2n+1) Equate to n+50 n(2n+1)=n+50 2n^2-50=0 n^2-25=0 n=+/-5 Since (2(-5)+1) choose 2=0 will invalidate the inequality, -5 is rejected. => n=5